# Bayesian estimation of cross-coupling and reflection systematics in 21cm array visibility data

### v1

@@ 5 Results
We present systematic modelling results in autocorrelations and cross-correlations for varying noise levels: high-noise, corresponding to a single time integration, reduced noise, which assumes there has been time-averaging of visibilities, and very-low noise, where the noise power spectrum is approximately $10\%$ that of the EoR. In this last case, plots of the signal loss metric as a function of delay are included, as we are able to mitigate the systematics down to, or close to, the true power spectrum. For the higher-noise cases, we can ideally only mitigate down to this noise level, so it is not as informative to include plots of the respective signal loss metrics. Where possible, we compare our results to that of Kern et al. (2019) when their methods are applied to essentially the same data. The only difference is due to the random nature of the noise generated, but the average overall power of the noise remains the same.
An issue encountered during sampling was that the foreground and EOR model parameters would attempt to fit both their associated components and the systematics as there is no inherent way for the model to distinguish between the power contributions from each component. This would result in the model systematics parameters underestimating the strength of the mock systematics, leading to excess power and residual systematics following systematics subtraction.
This was mainly an issue in regions of delay space containing foregrounds, EoR, and a large number of systematics, for example $\tau\lesssim 600\>\rm{ns}$ in the autocorrelations. Here, in combination with the EoR, non-negligible foreground power, and closely-spaced subreflections, the model struggled to obtain parameter values which would lead to complete mitigation of the systematics in noise-free scenarios. Instead, it would overestimate the contributions from the EoR and foregrounds, leading to residual systematics following mitigation. In delay ranges where foregrounds are not present, and which contain fewer classes of systematics, which themselves are better distributed in delay, mitigation performs better.
@@ 5.1 Autocorrelation Systematics
Fig. 8 shows the sampling results for the full set of systematics (described in Section 3) in a 10.73 second integration autocorrelation power spectrum using the priors and sampler settings described in Sections 4.3 and 4.4, respectively. For this plot, 30,000 power spectra are formed from the 30,000 samples, from which the $95\rm{th}$ percentile is calculated, denoted by the orange region.
Figure 8: Top: A comparison between the autocorrelation mock data (black solid line) and the model results (orange region), which sampled for two cable reflections, 20 subreflections, and 10 cross-coupling spectra in a 10.73 second integration. The mock power spectrum consists of the foreground signal, EoR signal, and systematics, with the noise (grey dotted line) being taken into account when sampling. Note that no constraints can be placed on features below the noise ($1000\lesssim\tau\lesssim 1200$). The model is only centred on the mock power spectrum in this region because of the priors. Being below the noise level simply results in the model filling the prior space for these particular features. The model was run for 30,000 sampling steps, and the 95th percentile of these samples is shown here. The true power spectrum is denoted by the green dashed line, and contains only the foreground and EoR components.
@@
In the autocorrelation analysis, the full set of cable reflections, subreflections, and cross-coupling spectra is added to the mock data. While cross-coupling is much weaker in the autocorrelations than in the cross-correlations, the power is still non-negligible. However, there was little to no constraint on the cross-coupling parameters in this case, which was expected as they are below the noise level.
@@
Figure 9: Using the model results in Fig. 8, an attempt is made to subtract the systematics from the autocorrelation mock + noise data. The red region corresponds to the 95% confidence interval of the 30,000 recovered power spectra. The blue dot-dashed line corresponds to the results of Kern et al. (2019) when their calibration is applied to the data.
@@
Using the above sampling results, the systematics are subtracted from the noisy mock autocorrelation data (i.e. the noise power spectrum is added to the mock power spectrum), resulting in the power spectrum of Fig. 9. A recovered power spectrum is formed for each of the 30,000 samples, and as before the $95\rm{th}$ percentile is calculated and shown as bands as a function of delay. We are able to mitigate the systematics down to the noise, which is the best that can be expected. Ideally, the objective in this case is for the band/confidence region to match the true power spectrum as closely as possible, where the true data only contains foregrounds, the EoR signal, and instrumental bandpass.
While our recovered percentile region is mostly centred on this noise level, note that random unit gaussians are used to form the noise power spectra. It is not expected that the percentile regions should be perfectly centred on the noise spectrum, as the precise features of the noise spectrum will differ between model runs, and between plots, despite the overall noise level being consistent.
@@
The calibration results of Kern et al. (2019) when applied to the mock data are also included in Fig. 9. We see essentially the same behaviour, i.e. the systematics being mitigated down to the noise level. Again, the disparities between both our results and the noise spectrum are due to the randomisation present in the noise.
@@
As it is done in Kern et al. (2019), we model the reflection systematics in the autocorrelations and subsequently apply the results to the corresponding cross-correlation. However, a cross-correlation visibility with a noise level consistent with the autocorrelations (i.e. derived from a single time integration, see Fig. 7) sees no useful results when the reflections are subtracted. When the results are applied, we effectively only recover the noisy cross-correlation, where the noise is on the order of $10^{8}\,\rm{Jy}^{2}\rm{Hz}^{2}$ in the power spectrum. The reflection systematics peak at approximately $10^{5}\,\rm{Jy}^{2}\rm{Hz}^{2}$. In this case, the noise is too strong to see any noticeable difference between the reflection corrupted and recovered power spectra, and so we do not present this particular result.
@@
We omit the results for reflection mitigation in a reduced-noise case, where the noise is lower than in the above results, but still above the EoR level. As before, the systematics can be mitigated down to the noise level in the autocorrelation. In the cross-correlations, the noise is, again, well-above the EoR power spectrum, so there is not much to be learned in this case.
@@
If, instead, the model is run on visibilities with very low noise, where the its power spectrum is $\sim 10\%$ the EoR’s level, we are then able to mitigate the systematics in the autocorrelations to a level closer to the true power spectrum (Fig. 10). There remains some residual systematics with powers of around one to two orders of magnitude above the true power spectrum. This is primarily due to there being no means for the model to differentiate between the contributions from the foregrounds, signal, and systematics in this one-dimensional regime. The excess power in Fig. 10 is due to the foregrounds and EoR models fitting a portion of the systematics power, leading to an underestimation of the systematics power, which is ultimately insufficient in completely mitigating the systematics.
@@
Kern et al. (2019)’s method mainly suffers from confusion, where it is difficult to differentiate between the contributions from individual systematic features themselves to the overall power. While our model may suffer from this as well, it does not appear to the main limiting factor. While we retain the same number of samples for Fig. 10, as well as the formation of the 95th percentile, the bands do become much narrower as a result of the minimal noise, leading to tighter constraints on the systematics.
Taking these results and applying them to the low-noise cross-correlation visibilities, we are mostly able to recover the true signal in Fig. 11 (top). With the reduction in the noise and the subsequent narrowing of the percentile bands, however, there are a number of delays where our results do not make a recovery. Nevertheless, performance in this regime is very good. We opt not to include the result of Kern et al. (2019)’s method for clarity, but the performance is comparable in this case. The bottom plot of Fig. 11, the signal loss metric, does provide a direct comparison between the two results, however. Our 95th percentile bands are not only very narrow (a result of the reduced noise), but adhere very closely to $R_{3}=1$, suggesting we are resilient to signal loss, much like Kern et al. (2019). Kern et al. (2019)’s method displays a large amount of excess power at $\tau\approx 900\,\rm{ns}$, but this is likely due to initial parameter choices rather than underlying performance. This is also where the first cross-coupling features in the autocorrelations appear, which would understandably limit the reflection calibration performance. Apart from this feature, the signal loss characteristics between the two results are very similar.
@@
Figure 10: The recovered autocorrelation power spectrum when applied to data with very low noise. As a result of the reduced noise, the width of the percentile bands becomes very narrow.
@@
Figure 11: Top: The recovered cross-correlation power spectrum (light-red band) when the results of Fig. 10 are used to subtract the reflections from the 14.7 m east-west cross-correlation mock data (grey dot-dashed line) in a low noise regime (denoted by the grey dotted line). We omit the results of Kern et al. (2019) for clarity and to prevent too many overlapping lines, but their model does perform similarly to ours. The objective, as before, is to recover the true power spectrum (black dashed line). The bottom plot, the signal loss metric, provides a comparison between our and Kern et al. (2019)’s results. The signal loss metric corresponding to our model is denoted by the light-red band, and that of Kern et al. (2019) by the blue dashed line. The $R_{3}(\tau)=1$ line is the brown, dot-dashed line.
@@ 5.2 Cross-coupling Systematics
Figure 12: The model results for cross-coupling systematics in a reduced noise 14.7 m east-west cross-correlation, where the integration time has been increased from 10.73 s to 1000 s. As before, the noise is plotted separately from the mock data for illustration purposes. The orange region is the 95th percentile of the model power spectra samples.
Figure 13: Using the results of Fig. 12, the cross-coupling results are subtracted from the mock + noise data. The 95th percentile of all recovered power spectra from our model is denoted by the red region. The recovered spectrum from Kern et al. (2019)’s SVD method is denoted by the blue dot-dashed line. The aim is for the recovered spectra to match the true, systematics-free power spectrum (green dashed line), but the noise (grey dotted line) limits this.
@@
Figure 14: Top: The recovered cross-correlation power spectrum (red region) in the very-low-noise regime (denoted by the grey dotted line). This is again compared to the SVD method of Kern et al. (2019) (blue dot-dashed line), and the true, systematics-free power spectrum (green dashed line). Bottom: The corresponding signal loss results, with ours being the red region, and Kern et al. (2019)’s being the blue dashed line. The brown, dot-dashed line denotes $R_{3}(\tau)=1$.
@@
This necessitates a reduction of the noise in order to constrain the cross-coupling systematics to any appreciable degree. Fig. 12 shows the results when the integration time is increased from 10.73 seconds to 1000 seconds, i.e. assuming that around 100 visibilities have been time averaged. While it is not so simple when considering observational data, we do not explore any complicating factors, such as the changing sky signal over time, and instead assume we can simply average as many visibilities as desired.
@@
In this case, we are able to strongly constrain the highest amplitude cross-coupling peaks, with the level of constraint reducing as the noise level is approached. The peaks below the noise level are not constrained at all, as is expected. Using these results, we can mitigate the cross-coupling systematics down to the noise level (Fig. 13), performing similarly to the SVD method of Kern et al. (2019). Again, the standard deviation used to produce the noise is identical between our result and Kern et al. (2019)’s, but the draws from the normal distributions in Eq. 15 differ due to the randomisation, so there is an expected level of disparity between these two results.
@@
An integration time of $1000\>\rm{seconds}$ is considerably higher than the overall time averaging cadence in HERA analysis, which is $214$ seconds. Furthermore, systematics calibration is carried out on data averaged to a $21.4$ second cadence, with the remaining coherent time averaging occurring after the systematics are removed (The HERA Collaboration et al., 2023). As such, this is not a representative example of noise reduction in the analysis pipeline, but nevertheless noise needs to be reduced in some manner in order to constrain cross-coupling systematics at all.
@@
The cross-coupling modelling performance was also tested in the very-low-noise regime, with the noise power spectrum being $\sim 10\%$ that of the EoR. This is a substitute for the noise-free case, as the HMC sampler requires a noise estimate for the likelihood function. Using noise levels this low slows sampling greatly, so it was opted not to reduce it further.
@@
We omit the comparison between the model and mock power spectra, and instead only show the recovered power spectra in Fig. 14 (top). Kern et al. (2019)’s SVD method, wherein a systematics model is formed from the mock data itself, performs well and is able to essentially remove the cross-coupling systematics completely. Our sampling method performs similarly, and is able to mitigate essentially the entire cross-coupling feature. To achieve this, however, a different model initialisation was required. Previous results all used, and performed best with the ADVI initialiser (discussed in Section 4.3). When modelling the cross-coupling in this low-noise cross-correlation, the systematics could only be mitigated by around two to four orders of magnitude, leaving a significant level of residual systematics. Opting instead for the jitter+adapt_diag initialiser improved this significantly. Here, a mass matrix for the HMC is formed, and a uniform jitter is applied to the starting sampler’s points. This is an overall simpler method of initialisation in comparison to ADVI, and it is not clear to us why performance is worse for the preferred initialiser.
The percentile bands of the recovered power spectra are narrowest at lower delays, corresponding to mitigation of the strongest cross-coupling peaks, which are easier to constrain. Despite the low noise, and that cross-coupling is the only class of systematic being modelled, the bands are still fairly broad towards higher delays, more so than the low-noise autocorrelation results. This is possibly due to the aforementioned confusion between individual coupling features, or the sampler being unable to distinguish between the power contributions from the coupling and EoR. Nevertheless, the model performs well in this regime.
In terms of signal loss, both our and Kern et al. (2019)’s SVD approach perform similarly. The fractions of signal loss and residual systematics are of similar levels. There are a handful of delays where our model fails to make a recovery of the true power spectrum within the 95th percentile, but the percentile bands are mostly centered on the true spectrum.
6 Summary

### v2

@@ 5 Results
We present systematic modelling results in autocorrelations and cross-correlations for varying noise levels: high-noise, corresponding to a single time integration, reduced noise, which assumes there has been some level of noise reduction prior to modelling, and very-low noise, where the noise power spectrum level is approximately $10\%$ that of the EoR. In this last case, plots of the signal loss metric as a function of delay are included, as we are able to mitigate the systematics down to, or close to, the true power spectrum. For the higher-noise cases, we can ideally only mitigate down to this noise level, so it is not as informative to include plots of the respective signal loss metrics. Where possible, we compare our results to that of Kern et al. (2019) when their methods are applied to the same data.
@@ 5.1 Autocorrelation Systematics
Fig. 7 shows the sampling results for the full set of systematics (described in Section 3) in a single 10.73 second integration autocorrelation power spectrum using the priors and sampler settings described in Sections 4.5 and 4.6, respectively. For this plot, 10,000 power spectra are formed from the 10,000 samples, from which the $95\rm{th}$ percentile is calculated, denoted by the orange region.
Figure 7: A comparison between the autocorrelation mock data (black solid line) and the model results (orange region), which sampled for two cable reflections, 20 subreflections, and 10 cross-coupling spectra in a single 10.73 second integration. The mock power spectrum consists of the foreground, signal, and systematics, with the noise (grey dotted line) being taken into account when sampling. No constraints can be placed on features below the noise ($1000\lesssim\tau\lesssim 1200$). The model is only centred on the mock power spectrum in this region because of the priors. Being below the noise level simply results in the model filling the prior space for these particular features. The model was run for 10,000 sampling steps, and the 95th percentile of these samples is shown here. The true power spectrum is denoted by the green dashed line, and contains only the foreground and EoR.
@@
In the autocorrelation analysis, the full set of cable reflections, subreflections, and cross-coupling spectra is added to the mock data. While cross-coupling is much weaker in the autocorrelations than in the cross-correlations, the power is still non-negligible. However, there was little to no constraint on the cross-coupling parameters in this case, which was expected as the majority of the peaks are below the noise level.
@@
Figure 8: After mitigating the systematics in thirty time samples, each separated by 21.46 seconds, the incoherent time average of the recovered spectra are found, whereafter the percentile regions are formed, i.e. we go from $30\times 10,000$ power spectra to $10,000$ mean spectra. The results of Kern et al. (2019) also correspond to the averaged recovered spectra when applied to the same data. The grey dotted line denotes the noise level for a single integration, and shows that noise can be mitigated following systematics removal.
@@
The model used above is applied to thirty visibilities each separated by 21.46 seconds. The systematics are mitigated from these visibilities, producing 10,000 recovered power spectra for each individual time sample, which are then incoherently averaged across time. From these 10,000 time-averaged recovered spectra, the 95th percentile is found, and is shown in Fig. 8. Here, the thirty mock and true power spectra are time averaged as well, and are plotted against the noise level for a single integration. We are able to mitigate the systematics down to the single integration noise level, and thereafter incoherently average to further reduce the noise in our recovered spectra. Ideally, the objective in this case is for the band/confidence region to match the true power spectrum as closely as possible, where the true data only contains foregrounds, the EoR signal, and instrumental bandpass.
@@
The calibration results of Kern et al. (2019) when applied to the mock data are also included in Fig. 8. Again, the calibration is applied to thirty time samples, and the power spectra are incoherently averaged across time. We see essentially the same behaviour, i.e. the systematics being mitigated down to the single-integration noise level, with there being further reduction in the noise after time averaging.
@@
As it is done in Kern et al. (2019), we model the reflection systematics in the autocorrelation and subsequently apply the results to the corresponding cross-correlation. However, a cross-correlation visibility with a noise level consistent with the autocorrelations (i.e. derived from a single time integration, see Fig. 6) sees no useful results when the reflections are subtracted. When the results are applied, we effectively only recover the noisy cross-correlation, where the noise is on the order of $10^{8}\,\rm{Jy}^{2}\rm{Hz}^{2}$ in the power spectrum. The reflection systematics peak at approximately $10^{5}\,\rm{Jy}^{2}\rm{Hz}^{2}$. In this case, the noise is too strong to see any noticeable difference between the reflection corrupted and recovered power spectra, and so we do not present this particular result.
@@
We omit the results for reflection mitigation in a reduced-noise case, where the noise is lower than in the above results, but still above the EoR level. As before, the systematics can be mitigated down to the noise level in the autocorrelation. In the cross-correlations, the noise is, again, well-above both the EoR power spectrum and the reflection systematics, so there is not much to be learned in this case.
@@
If, instead, the model is run on visibilities with very low noise, where its power spectrum is $\sim 10\%$ the EoR’s level, we are then able to mitigate the systematics in the autocorrelations to a level closer to the true power spectrum (Fig. 9). We mainly test this case to assess the signal loss properties of our model, and whether the mitigation removes primarily the systematics as it should, or if it suppresses the signal to too great a degree. There remains some residual systematics with powers of up to three orders of magnitude above the true power spectrum. This is primarily due to confusion between the systematics. Kern et al. (2019)’s method mainly suffers from this as well, where it is difficult to differentiate between the contributions from individual systematic features to the overall power.
@@
While we retain the same number of samples for Fig. 9, as well as the formation of the 95th percentile, the bands do become much narrower as a result of the minimal noise, leading to tighter constraints on the systematics.
By taking these results for the systematics parameters and applying them to the low-noise cross-correlation visibilities, we are mostly able to recover the true signal in Fig. 10 (top). With the reduction in the noise and the subsequent narrowing of the percentile bands, however, there are a number of delays where our results make somewhat poorer recovery, i.e. at $\sim 750$ ns and $\sim 900-1100$ ns. This is primarily a result of the performance in the corresponding autocorrelation (Fig. 9) which carries over to the cross-correlation systematic mitigation. As detailed previously, the poor recovery in the autocorrelation is a result of confusion between the closely-spaced systematics. The clearest example of this is at $\sim 900-1100$ ns in Fig. 9. Here, both confused subreflections and confused coupling systematics occupy the same delay range, leading to the largest amount of residual systematics, $\sim 3$ orders of magnitude above the true power spectrum. This results in poorer recovery of the corresponding cross-correlation at similar delays, i.e. residual reflections or oversubtraction in Fig. 10.
Where there are fewer classes of overlapping systematics, performance is improved, for example at $200-900$ ns. While the model is still confusion-limited, the effects are not as pronounced when applied to the cross-correlation. Overall, however, this effect of confusion is ultimately a result of being unable to efficiently explore the parameter space. Normally one could choose wider, less-specific priors at the cost of slower sampling speed. However, sampling speeds already suffer as a result of the low noise, which produces prohibitive run-times. The alternative approach to mitigate confusion would be to choose extremely specific priors, but that would imply very good knowledge of the parameters and would defeat the purpose of sampling. Were the properties of the systematics known to such a degree, then a non-statistical mitigation approach would be sufficient.
Nevertheless, the model is capable of making a good recovery of reflection-corrupted cross-correlations. We opt not to include the result of Kern et al. (2019)’s method for clarity, but the performance is comparable in this case. The bottom plot of Fig. 10, the signal loss metric, does provide a direct comparison between the two results, however. For the most part, our 95th percentile band adheres very closely to $R_{3}=1$, suggesting we are resilient to signal loss, much like Kern et al. (2019). We see noticeably more residual systematics and signal loss at $900\lesssim\tau\lesssim 1100$ ns, a result of the overlapping reflections and coupling in the autocorrelation, but overall, the signal loss characteristics between the two results are very similar.
@@
Figure 9: The recovered autocorrelation power spectrum when applied to data with very low noise. Here, we simply set the noise level to approximately 10% the power of the EoR, rather than assume any particular integration time. As a result of the reduced noise, the width of the percentile bands becomes very narrow.
@@
Figure 10: Top: The recovered cross-correlation power spectrum (light-red band) when the results of Fig. 9 are used to subtract the reflections from the 14.7 m east-west cross-correlation mock data (grey dot-dashed line) in a low noise regime (denoted by the grey dotted line). We omit the results of Kern et al. (2019) for clarity and to prevent too many overlapping lines, but their model does perform similarly to ours, and their signal loss result is shown in the bottom plot. The objective, as before, is to recover the true power spectrum (black dashed line). The bottom plot, the signal loss metric, provides a comparison between our and Kern et al. (2019)’s results. The signal loss metric corresponding to our model is denoted by the light-red band, and that of Kern et al. (2019) by the blue dashed line. The $R_{3}(\tau)=1$ line is the brown, dot-dashed line, and denotes perfect recovery.
@@ 5.2 Cross-coupling Systematics
Figure 11: The model results for cross-coupling systematics in a reduced noise 14.7 m east-west cross-correlation, where the integration time has been increased from 10.73 s to 1073 s, i.e. we have not incoherently averaged multiple power spectra here, but simply reduced the noise prior to modelling, for testing purposes. As before, the noise is plotted separately from the mock data for illustration purposes. The orange region is the 95th percentile of the 10,000 model power spectra samples. As discussed in Section 4.2, we omit a signal component, which is why the power spectrum is unmodelled at delays away from the foreground and systematics peaks.
Figure 12: The power spectrum recovery when jointly modelling thirty time integrations separated by 21.46 seconds, recovering the underlying power spectra, and thereafter incoherently time averaging. Here, 10,000 samples are taken. The results of Kern et al. (2019), the true power spectra, and the mock power spectra have all been averaged as well. The noise power spectrum corresponds to the level prior to incoherent averaging, and is the power at which sampling takes place.
Figure 13: An example of the log posterior density during the tuning (burn-in) phase, corresponding to the sampling results of Fig. 12. This plot shows the first 250 of 5,000 tuning steps for the two chains, demonstrating that the model tunes fairly quickly for this particular result. Sampling commences after this stage.
@@
Figure 14: Top: The recovered cross-correlation power spectrum (red region) in the very-low-noise regime (denoted by the grey dotted line), where we assume the noise is below the signal. This is for a single mock spectrum, and no incoherent averaging is carried out for this result. This is again compared to the SVD method of Kern et al. (2019) (blue dot-dashed line), and the true, systematics-free power spectrum (green dashed line). Bottom: The corresponding signal loss results, with ours being the red region, and Kern et al. (2019)’s being the blue dashed line. The brown, dot-dashed line denotes $R_{3}(\tau)=1$ for perfect recovery.
@@
This necessitates a reduction of the noise in order to constrain the cross-coupling systematics to any appreciable degree. Fig. 11 shows the results when the integration time in Eq. 15 is increased from 10.73 seconds to 1073 seconds, i.e. assuming that noise has been reduced prior to systematics mitigation. While it is not so simple when considering observational data, we do not explore any complicating factors, such as the changing sky signal over time, and instead assume a reduced noise level so that we can test our model for this systematic.
@@
In this case, we are able to strongly constrain the highest amplitude cross-coupling peaks, with the level of constraint reducing as the noise level is approached (Fig. 11). The peaks below the noise level are not constrained at all, as is expected. Using these results, we can mitigate the cross-coupling systematics down to the noise level in a single integration, performing similarly to the method of Kern et al. (2019) when the first 15 SVD modes are used.
@@
An integration time of $1073\>\rm{seconds}$ is considerably higher than the overall time averaging cadence in HERA analysis, which is $214$ seconds. Furthermore, systematics calibration is carried out on data averaged to a $21.4$ second cadence, with the remaining coherent time averaging occurring after the systematics are removed (The HERA Collaboration et al., 2023). As such, this is not a representative example of noise reduction in the analysis pipeline, but nevertheless noise needs to be reduced in some manner in order to constrain cross-coupling systematics at all.
As with the reflection modelling, we apply the coupling model to thirty time integrations, mitigate this systematic, and incoherently time average all of the samples. However, following the same process as we did for the autocorrelations produced poor results. Namely, sampling each integration individually, recovering the power spectra, and thereafter averaging, did not produce a further reduction of the noise in the delay range of the systematics. This suggests that the model could constrain the coupling systematics only well enough to mitigate them down to the noise level ($\sim 10^{7}\,\rm{Jy}^{2}\rm{Hz}^{2}$), but left residual systematics hidden by this noise. Attempting to average the recovered spectra made this additional structure evident, and the incoherently averaged power spectrum remained at $\sim 10^{7}\,\rm{Jy}^{2}\rm{Hz}^{2}$ at delays of $900\lesssim\tau\lesssim 1100\,\rm{ns}$, while delays dominated by noise saw a decrease in power.
If, instead, we jointly model all thirty time integrations simultaneously, then the results are improved. Here, we have thirty terms in our likelihood function, and the systematics priors and parameters are shared between all of them. In other words, all integrations have the same systematics solution solved for at once, rather than solving for each visibility individually, which is what was described in the paragraph above. Using the result from the joint model, we subtract the systematics from the thirty visibilities, and incoherently average the power spectra.
This is compared to the time averaged mock and true power spectra, as well as the recovered spectra from the Kern et al. (2019) calibration in Fig. 12. We can mitigate the systematics down to the noise level for all thirty power spectra, and incoherently averaging produces a further reduction in the noise. Our result is comparable to the result of Kern et al. (2019). Outside of differences in the finer features, the incoherently averaged recovered power spectra are at a power of $\sim 10^{5}\,\rm{Jy}^{2}\rm{Hz}^{2}$, a reduction of around two orders of magnitude from the initial noise level. Our result has the benefit of an associated uncertainty, and for the most part there is agreement between the two results within this uncertainty.
We also provide an example of the tuning/burn-in phase for this particular result. Pymc3 tracks the log posterior density, which is proportional to the product of the likelihood and prior of the model (Lynch, 2007). The aim is for the model to maximise the posterior density. As expected, there is an increase in this quantity as the model explores, and eventually settles in, the posterior space. There are 5,000 tuning steps for both chains, and following this the model draws a total of 10,000 samples, corresponding to the results of Fig. 12.
@@
The cross-coupling modelling performance was also tested in the very-low-noise regime, with the noise power spectrum being $\sim 10\%$ that of the EoR. Using noise levels this low slows sampling severely, so it was opted not to reduce it further, and only a single integration was modelled.
@@
We omit the comparison between the model and mock power spectra, and instead only show the recovered power spectra in Fig. 14 (top). Kern et al. (2019)’s SVD method performs well and is able to essentially remove the cross-coupling systematics completely. Our sampling method performs similarly, although our model results in a fair amount of residual systematics, while Kern et al. (2019)’s is more prone to oversubtraction. To produce our result, a different model initialisation was required. Previous results all performed best with the ADVI initialiser (discussed in Section 4.5). When modelling the cross-coupling in this low-noise cross-correlation, the systematics could only be mitigated by around two to four orders of magnitude, leaving a significant level of residual systematics. Opting instead for the jitter+adapt_diag initialiser improved this significantly. Here, a mass matrix for the HMC is formed, and a uniform jitter is applied to the starting sampler’s points. This is an overall simpler method of initialisation in comparison to ADVI, and it is not clear to us why performance is worse for the preferred initialiser.
The mitigation of the strongest cross-coupling peaks appears somewhat easier for our model, given their strength and relative lack of confusion. The weaker, more confused peaks show poorer mitigation, resulting in these residual systematics. Nevertheless, the model performs well in this regime.
In terms of signal loss, both our and Kern et al. (2019)’s SVD approach perform similarly in terms of magnitude, although as mentioned, our method has a predisposition to incomplete mitigation, while Kern et al. (2019)’s tends to oversubtract more frequently.
5.3 Alternate/Incorrect Models
We briefly discuss the results when the model for the mock visibilities is incorrect. Our fiducial mock data consists of 20 subreflections and 10 coupling peaks. We evaluate the results when:
  1. 1.
    Our fitting model contains too few or too many of these systematic features. Specifically, we test the cases of there being either 17 or 23 subreflections, and the cases of there being either 8 or 12 coupling peaks in our model.
  2. 2.
    There is the correct number of systematics features, but the parameters for them have been randomised to a degree, i.e. little effort has been put into making accurate initial estimates.
Overall, the results are fairly consistent regardless of whether too many, too few, or randomised systematics features are assumed. All cases result in an excess of power following systematics mitigation, rather than an oversubtraction. Estimates of the systematics which are ‘near enough’ to the true values result in slight suppression, while very poor estimates result in an addition of power. This is a result of the exponential terms in the systematics equations, Eqs. 6 and 10, which themselves contain phases. For poor delay estimates, in particular, subtraction of model systematics does not lead to suppression of power, but rather addition.
Figure 15: Examples of the mean recovered power spectra when the assumed systematics model differs from that present in the mock data. The fiducial coupling model contains ten peaks. The top plot shows the incoherently averaged recovered cross-correlation power spectra when only eight peaks are assumed, and the middle plot when twelve are assumed. The bottom plot shows the averaged recovered autocorrelations when the correct number of subreflections are assumed, but the initial delay estimates are randomly chosen. Similar results are observed across the different scenarios not shown here, i.e. too many/too few/randomised systematics in either the autocorrelations or cross-correlations.
Both reflection modelling in the autocorrelations and coupling modelling in the cross-correlations show the same general behaviour for scenario (i). When there are too few systematics features, e.g. Fig. 15 (top), the recovered power spectra exhibit excess power at the delays of the unmodelled systematics ($\sim 1100-1200$ ns). When there are too many assumed systematics, this produces similar excesses in power. Fig. 15 (middle) shows an excess of power in the recovered spectra at delays where two too many coupling peaks are introduced ($\sim 900-1000$ ns).
Assuming the correct number of systematics in the correct delay range, but randomly placing the initial systematics peaks (with wider priors to account for this) produces somewhat more reasonable results, but still results in insufficient recovery. For most delays, and in the autocorrelations, recovery can be made down to the noise level for single time integrations when the subreflection peaks are randomly placed between $\sim 200-1000$ ns. However, attempting to incoherently average the recovered power spectra, as in Fig. 15 (bottom), does not produce any meaningful level of noise reduction, suggesting there are significant amounts of residual subreflection systematics.
With randomised initial systematics parameters in the negligible noise scenario (not shown here), we see little to no mitigation of the systematics. This is mainly due to the low level of the noise, which results in the sampler less efficiently exploring the posterior space. Nevertheless, as before, we do not see any oversubtraction or signal loss, but this does suggest that a fair amount of effort should be put into setting the priors. This is not unlike the requirements of Kern et al. (2019)’s reflection mitigation method, which requires estimates of the reflection peak delay. However, Kern et al. (2019)’s coupling mitigation does have an advantage over ours, as only the delay range of the coupling peaks is required, whereas we again need initial delay, amplitude, and phase estimates.
Generally, poor priors for the systematics in our model are unlikely to result in oversubtraction, and are instead prone to induce excess power in the recovered spectra. It is not immediately obvious how to distinguish from the recovered spectra if there are too many assumed systematics features in the model versus too few, but it was noticed that when too few were assumed, the excess power demonstrated more of a typical peak shape. Assuming too many features resulted in wider confidence intervals in the region of the additional features.
6 Discussion